# Rates of convergence in long time asymptotics of an alignment model with symmetry breaking

### v2

@@ 5.2. A coercivity result
$|u_{f_{\mathrm{ini}}}-u_{*}|$ $=\left|\int_{\mathbb{R}^{d}}v\,g_{\mathrm{ini}}\,G_{*}\,dv\right|\leq\frac{W_{2}^{\frac{1}{2}}}{\eta(D)}\,Q_{1,u_{*}}[g_{\mathrm{ini}}]<\delta.$
For all $t\geq 0$ such that $|u_{f}(t)-u_{*}(t)|<\delta$, using Lemma˜5.7 and (54), we have:
$\frac{d}{dt}Q_{1,u_{*}}[g(t,.)]$ = $=-\,2Q_{2,u_{*}}[g(t,.)]+2R[g(t,.)]$
 $\leq-\,2Q_{2,u_{*}}[g]+2K_{D}Q_{2,u_{*}}[g]Q_{1,u_{*}}[g]^{\frac{1}{2}}\leq-\,2Q_{2,u_{*}}[g]\left(1-K_{D}Q_{1,u_{*}}[g]^{\frac{1}{2}}\right).$
Since $Q_{1,u_{*}}[g_{\mathrm{ini}}]<A_{D}\leq\frac{1}{K_{D}^{2}}$, we obtain that $Q_{1,u_{*}}[g]$ is nonincreasing hence, for $t\geq 0$,
$|u_{f}(t)-u_{*}(t)|$ $=\left|\int_{\mathbb{R}^{d}}vgG_{u}dv\right|\leq\frac{W_{2}^{\frac{1}{2}}}{\eta(D)}Q_{1,u_{*}}[g]\leq\frac{W_{2}^{\frac{1}{2}}}{\eta(D)}Q_{1,u_{*}}[g_{\mathrm{ini}}]<\delta.$
We conclude that $u_{f}(t)\in\mathcal{N}$ for all time $t\geq 0$. Hence, for all $t\geq 0$ we have
$\frac{d}{dt}Q_{1,u_{*}}[g(t,.)]\leq-\,2\beta(D)^{2}\,Q_{1,u_{*}}[g(t,.)]\left(1-K_{D}Q_{1,u_{*}}[g(t,.)]^{\frac{1}{2}}\right),$
@@
$Q_{1,u_{*}}[g]\leq\frac{Q_{1,u_{*}}[g_{\mathrm{ini}}]}{\left(1-K_{D}\,Q_{1,u_{*}}[g_{\mathrm{ini}}]^{1/2}\right)^{2}}\,e^{-2\beta(D)^{2}t}\,\,\,\forall t\geq 0\qed$
@@
$\lim_{t\rightarrow\infty}D|\,v_{g}(t)|=\lim_{t\rightarrow\infty}\left|u_{f}(t)-u_{*}(t)\right|=\lim_{t\rightarrow\infty}\mathrm{dist}(u_{f},\mathcal{S})=0.$ (56)
For $t$ big enough, $|u_{f}(t)|\geq r(D)/2$ and we know from (31) that $\lim_{t\rightarrow}u_{f}^{\prime}(t)=0$. Hence, using the inequality $|u_{*}^{\prime}(t)|\leq 2{|u_{f}^{\prime}|}/{|u_{f}|}$, we obtain by Lemma˜2.2
$\lim_{t\rightarrow\infty}|u_{*}^{\prime}(t)|=0.$ (57)
Thanks to (56), (57) and (53), we conclude that for all $\epsilon\in(0,1)$, there exists a time $t_{0}>0$ such that for all $t\geq t_{0}$, we have $R[g(t,.)]\leq\epsilon\,Q_{2,u_{*}}[g(t,.)]$, and
$\frac{d}{dt}Q_{1,u_{*}}[g(t,.)]$ = $=-\,2Q_{2,u_{*}}[g(t,.)]+2R[g(t,.)]$
 $\leq-\,2(1-\epsilon)Q_{2,u_{*}}[g(t,.)]\leq-\,2\beta(D)^{2}(1-\epsilon)Q_{1,u_{*}}[g(t,.)].$
As a consequence, $Q_{1,u_{*}}[g]$ converges to 0. Hence for $t_{0}$ big enough, we can apply Proposition˜5.2 to $Q_{1,u_{*}(t_{0})}[g(t_{0}+.)]$ and we obtain
${Q_{1,u_{*}}[g(t)]}\leq\frac{Q_{1,u_{*}}[g(t_{0})]}{\left(1-K_{D}\,Q_{1,u_{*}}[g{(t_{0})}]^{\frac{1}{2}}\right)^{2}}\,e^{-2\beta(D)^{2}(t-t_{0})}\text{ for all }t\geq t_{0}.$
@@
$f\log\left(\frac{f}{G_{u}}\right)-(f-G_{u})\leq\frac{(f-G_{u})^{2}}{G_{u}},$
@@
$\mathcal{H}[f(t,.)|G_{*}(t)]\leq\int_{\mathbb{R}^{d}}|f-G_{*}|^{2}G_{*}^{-1}dv=\lVert g\rVert_{2,u_{*}}^{2}\leq\frac{1}{\eta(D)^{2}}Q_{1,u_{*}}[g(t,.)].$ (58)
Let $\epsilon>0$ and $\mathcal{N}$ a neighborhood of $\mathcal{S}$ where (19) applies. Using Lemma˜2.2, there exists a time $t_{0}>0$, such that $u_{f}(t)\in\mathcal{N}$ for all $t\geq t_{0}$. Using (13), (15) and (58) we have for $t\geq t_{0}$:
$|u_{f}(t)-u_{*}(t)|^{2}\leq\frac{2}{\mu}(\mathcal{V}(u_{f})-\mathcal{V}_{*})\leq\frac{2}{\mu}(\mathcal{F}[f]-\mathcal{F}_{*})\leq\frac{2D}{\mu}\mathcal{H}[f|G_{u_{*}}]\leq\frac{2D}{\mu\eta(D)^{2}}Q_{1,u_{*}}[g(t,.)].$ (59)
@@
$|u_{*}(t)-u_{\infty}|^{2}$ $\leq\left|\int_{t}^{\infty}u_{*}^{\prime}(s)ds\right|^{2}\leq\frac{2\alpha W_{6}^{\frac{1}{2}}}{\beta(D)^{2}r(D)\eta(D)}\frac{Q_{1,u_{*}}[g(t_{0})]}{\left(1-K_{D}Q_{1,u_{*}}[g(t_{0})]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}(t-t_{0})}.$ (60)
@@
$|u_{f}(t)-u_{\infty}|^{2}$ $\leq 2\left(|u_{f}-u_{*}|^{2}+|u_{*}-u_{\infty}|^{2}\right)$
$\leq 2\left(\frac{2D}{\mu\eta(D)^{2}}+\frac{2\alpha W_{6}^{\frac{1}{2}}}{\beta(D)^{2}r(D)\eta(D)}\right)\frac{Q_{1,u_{*}}[g(t_{0})]}{\left(1-K_{D}Q_{1,u_{*}}[g(t_{0})]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}(t-t_{0})}.$ (61)
After taking into account (15) with $u=u_{\infty}$, (5.5) and (59), we obtain for $t\geq t_{0}$,
$\mathcal{H}[f(t,.)|G_{u_{\infty}}]$ $\leq\frac{1}{D}\left(\mathcal{F}[f]-\mathcal{F}_{*}+\frac{1}{2}|u_{f}-u_{\infty}|^{2}\right)$
$\leq\left(\frac{1}{\eta(D)^{2}}+\left(\frac{2}{\mu\eta(D)^{2}}+\frac{2\alpha W_{6}^{\frac{1}{2}}}{D\beta(D)^{2}r(D)\eta(D)}\right)\right)\frac{Q_{1,u_{*}}[g(t_{0})]}{\left(1-K_{D}Q_{1,u_{*}}[g(t_{0})]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}(t-t_{0})}.$
@@
Acknowledgements. This work has been supported by the Project Conviviality (ANR-23-CE40-0003) of the French National Research Agency.
© 2025 by the authors. This paper may be reproduced, in its entirety, for non-commercial purposes.

### v3

@@ 5.2. A coercivity result
$|\mathbf{u}_{f_{\mathrm{in}}}-\mathbf{u}_{*}|$ $=\left|\int_{\mathbb{R}^{d}}v\,g_{\mathrm{in}}\,G_{*}\,dv\right|\leq\frac{W_{2}^{\frac{1}{2}}}{\eta(D)}\,Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]<\delta.$
For all $t\geq 0$ such that $|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)|<\delta$, using Lemma˜5.7 and (55), we have:
$\frac{d}{dt}Q_{1,\mathbf{u}_{*}}[g(t,.)]$ = $=-\,2Q_{2,\mathbf{u}_{*}}[g(t,.)]+2R[g(t,.)]$
 $\leq-\,2Q_{2,\mathbf{u}_{*}}[g]+2K_{D}Q_{2,\mathbf{u}_{*}}[g]Q_{1,\mathbf{u}_{*}}[g]^{\frac{1}{2}}\leq-\,2Q_{2,\mathbf{u}_{*}}[g]\left(1-K_{D}Q_{1,\mathbf{u}_{*}}[g]^{\frac{1}{2}}\right).$
Since $Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]<A_{D}\leq\frac{1}{K_{D}^{2}}$, we obtain that $Q_{1,\mathbf{u}_{*}}[g]$ is nonincreasing hence, for $t\geq 0$,
$|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)|$ $=\left|\int_{\mathbb{R}^{d}}vgG_{\mathbf{u}}dv\right|\leq\frac{W_{2}^{\frac{1}{2}}}{\eta(D)}Q_{1,\mathbf{u}_{*}}[g]\leq\frac{W_{2}^{\frac{1}{2}}}{\eta(D)}Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]<\delta.$
We conclude that $\mathbf{u}_{f}(t)\in\mathcal{N}$ for all time $t\geq 0$. Hence, for all $t\geq 0$ we have
$\frac{d}{dt}Q_{1,\mathbf{u}_{*}}[g(t,.)]\leq-\,2\beta(D)^{2}\,Q_{1,\mathbf{u}_{*}}[g(t,.)]\left(1-K_{D}Q_{1,\mathbf{u}_{*}}[g(t,.)]^{\frac{1}{2}}\right),$
@@
$Q_{1,\mathbf{u}_{*}}[g]\leq\frac{Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]}{\left(1-K_{D}\,Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]^{1/2}\right)^{2}}\,e^{-2\beta(D)^{2}t}\,\,\,\forall t\geq 0\qed$
@@
$\lim_{t\rightarrow\infty}D|\,\mathbf{v}_{g}(t)|=\lim_{t\rightarrow\infty}\left|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)\right|=\lim_{t\rightarrow\infty}\mathrm{dist}(\mathbf{u}_{f},\mathcal{S})=0.$ (57)
For $t$ big enough, $|\mathbf{u}_{f}(t)|\geq r(D)/2$ and we know from (32) that $\lim_{t\rightarrow}\mathbf{u}_{f}^{\prime}(t)=0$. Hence, using the inequality $|\mathbf{u}_{*}^{\prime}(t)|\leq 2{|\mathbf{u}_{f}^{\prime}|}/{|\mathbf{u}_{f}|}$, we obtain by Lemma˜2.2
$\lim_{t\rightarrow\infty}|\mathbf{u}_{*}^{\prime}(t)|=0.$ (58)
Thanks to (57), (58) and (54), we conclude that for all $\epsilon\in(0,1)$, there exists a time $t_{0}>0$ such that for all $t\geq t_{0}$, we have $R[g(t,.)]\leq\epsilon\,Q_{2,\mathbf{u}_{*}}[g(t,.)]$, and
$\frac{d}{dt}Q_{1,\mathbf{u}_{*}}[g(t,.)]$ = $=-\,2Q_{2,\mathbf{u}_{*}}[g(t,.)]+2R[g(t,.)]$
 $\leq-\,2(1-\epsilon)Q_{2,\mathbf{u}_{*}}[g(t,.)]\leq-\,2\beta(D)^{2}(1-\epsilon)Q_{1,\mathbf{u}_{*}}[g(t,.)].$
As a consequence, $Q_{1,\mathbf{u}_{*}}[g]$ converges to 0. Hence for $t_{0}$ big enough, we can apply Proposition˜5.2 to $Q_{1,\mathbf{u}_{*}(t_{0})}[g(t_{0}+.)]$ and we obtain
${Q_{1,\mathbf{u}_{*}}[g(t)]}\leq\frac{Q_{1,\mathbf{u}_{*}}[g(t_{0})]}{\left(1-K_{D}\,Q_{1,\mathbf{u}_{*}}[g{(t_{0})}]^{\frac{1}{2}}\right)^{2}}\,e^{-2\beta(D)^{2}(t-t_{0})}\text{ for all }t\geq t_{0}.$
@@
$f\log\left(\frac{f}{G_{\mathbf{u}}}\right)-(f-G_{\mathbf{u}})\leq\frac{(f-G_{\mathbf{u}})^{2}}{G_{\mathbf{u}}},$
@@
$\mathcal{H}[f(t,.)|G_{*}(t)]\leq\int_{\mathbb{R}^{d}}|f-G_{*}|^{2}G_{*}^{-1}dv=\lVert g\rVert_{2,\mathbf{u}_{*}}^{2}\leq\frac{1}{\eta(D)^{2}}Q_{1,\mathbf{u}_{*}}[g(t,.)].$ (59)
Let $\epsilon>0$ and $\mathcal{N}$ a neighborhood of $\mathcal{S}$ where (19) applies. Using Lemma˜2.2, there exists a time $t_{0}>0$, such that $\mathbf{u}_{f}(t)\in\mathcal{N}$ for all $t\geq t_{0}$. Using (13), (15) and (59) we have for $t\geq t_{0}$:
$|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)|^{2}\leq\frac{2}{\mu}(\mathcal{V}(\mathbf{u}_{f})-\mathcal{V}_{*})\leq\frac{2}{\mu}(\mathcal{F}[f]-\mathcal{F}_{*})\leq\frac{2D}{\mu}\mathcal{H}[f|G_{\mathbf{u}_{*}}]\leq\frac{2D}{\mu\eta(D)^{2}}Q_{1,\mathbf{u}_{*}}[g(t,.)].$ (60)
@@
$|\mathbf{u}_{*}(t)-\mathbf{u}_{\infty}|^{2}$ $\leq\left|\int_{t}^{\infty}\mathbf{u}_{*}^{\prime}(s)ds\right|^{2}\leq\frac{2\alpha W_{6}^{\frac{1}{2}}}{\beta(D)^{2}r(D)\eta(D)}\frac{Q_{1,\mathbf{u}_{*}}[g(t_{0})]}{\left(1-K_{D}Q_{1,\mathbf{u}_{*}}[g(t_{0})]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}(t-t_{0})}.$ (61)
@@
$|\mathbf{u}_{f}(t)-\mathbf{u}_{\infty}|^{2}$ $\leq 2\left(|\mathbf{u}_{f}-\mathbf{u}_{*}|^{2}+|\mathbf{u}_{*}-\mathbf{u}_{\infty}|^{2}\right)$
$\leq 2\left(\frac{2D}{\mu\eta(D)^{2}}+\frac{2\alpha W_{6}^{\frac{1}{2}}}{\beta(D)^{2}r(D)\eta(D)}\right)\frac{Q_{1,\mathbf{u}_{*}}[g(t_{0})]}{\left(1-K_{D}Q_{1,\mathbf{u}_{*}}[g(t_{0})]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}(t-t_{0})}.$ (62)
After taking into account (15) with $\mathbf{u}=\mathbf{u}_{\infty}$, (5.5) and (60), we obtain for $t\geq t_{0}$,
$\mathcal{H}[f(t,.)|G_{\mathbf{u}_{\infty}}]$ $\leq\frac{1}{D}\left(\mathcal{F}[f]-\mathcal{F}_{*}+\frac{1}{2}|\mathbf{u}_{f}-\mathbf{u}_{\infty}|^{2}\right)$
$\leq\left(\frac{1}{\eta(D)^{2}}+\left(\frac{2}{\mu\eta(D)^{2}}+\frac{2\alpha W_{6}^{\frac{1}{2}}}{D\beta(D)^{2}r(D)\eta(D)}\right)\right)\frac{Q_{1,\mathbf{u}_{*}}[g(t_{0})]}{\left(1-K_{D}Q_{1,\mathbf{u}_{*}}[g(t_{0})]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}(t-t_{0})}.$
@@
Acknowledgements. This work has been supported by the Project Conviviality (ANR-23-CE40-0003) of the French National Research Agency. The author would like to thank Jean Dolbeault and Amic Frouvelle for their guidance and helpful discussions throughout this research, as well as the anonymous referee for valuable comments and suggestions.
© 2026 by the authors. This paper may be reproduced, in its entirety, for non-commercial purposes.